OpenStudy (anonymous):

find dy/dx by implicit differentiation. x^3+y^3=8

OpenStudy (anonymous):

@satellite73 please help

OpenStudy (anonymous):

\[x^{3}+y^{3}=8\] \[3x^{2}+3y^{2}\frac{ dy }{ dx }=0\] \[3y^{2}\frac{ dy }{ dx }=-3x^{2}\] \[\frac{ dy }{ dx }=-\frac{ x^{2} }{ y^{2} }\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

What is it?